Could you tell me how to use prime factorization becausem i do not understand how to do it like say for example like 525 how could i turn that in to a prime factorization
Answer by Earlsdon(6294)

Prime factorization simply means that you are to find all of the factors of the given number and each of these factors is a prime number. Remember that a prime number is divisble only by 1 and the number itself. Or, in other words, a prime number has only two factors, 1 and the number itself, whereas, composite (non-prime) numbers have three or more factors.
When you are doing these kinds of problems, it really helps to know the divisibilty rules.
Let's look at 525 and you can see by inspecton that it is divisible by 3 and by 5. If 525 is divisible by 3, then 3 is a factor of 525.
525/3 = 175: 3 is a prime number, and 175 is divisible by 5.
175/5 = 35: 5 is a prime number, and 35 is divisible by 5.
35/5 = 7: 5 is a prime number and so is 7
7/7 = 1
The prime factors of 525 are: 3, 5, 5, 7
You can check this by multipying these factors to see if you get 525. I'll leave this as an exercise for you.